OpenStudy (anonymous):

square root 75x^2 can anyone answer that?

OpenStudy (anonymous):

\[5x \sqrt3\]

OpenStudy (anonymous):

how dhid yu get that?

OpenStudy (anonymous):

note that 75 x^2 = 3 * 25 x^2 = 3*(5x)^2

jimthompson5910 (jim_thompson5910):

\[\sqrt{75x^2}=\sqrt{25*3x^2}=\sqrt{25}\sqrt{3}\sqrt{x^2}=5x\sqrt{3}\]

OpenStudy (anonymous):

yea i don't get that lol,, i learn better when sumone is actually showing me in person step by step i'll jst get a tutor noo big deal

jimthompson5910 (jim_thompson5910):

Basically I'm factoring 75 into factors where one factor is a perfect square. When I break up the square root, I can take the square root of 25 to get 5.

OpenStudy (anonymous):

oh ok i guess,,thanks